# Fixed-Time Input-to-State Stability for Singularly Perturbed Systems via Composite Lyapunov Functions

### v1

@@ 3.2 A Stylized Example
$\dfrac{\partial V}{\partial x}f(x,h(x),u)$ $=-2|x|^{r_{1}+1}-2|x|^{r_{2}+1}+2xu_{1}$
$\leq-V^{\frac{1}{2}(r_{1}+1)}-V^{\frac{1}{2}(r_{2}+1)}+|u|^{2},$ (20)
which shows that Assumption 2 is satisfied. To proceed with the boundary layer dynamics, we first establish the following useful lemma:
Lemma 3.4.
Given $\alpha>0$ and $x,u\in\mathbb{R}$ the following holds for $|x|>2|u|$:
$x\lceil x+u\rfloor^{\alpha}>2^{-\alpha}|x|^{\alpha+1}.$
Proof 3.5.
First assume $x<0$, which implies $|u|<-\frac{1}{2}x$. Since $\lceil x\rfloor^{\alpha}$ is strictly increasing in $x$, we have
$\lceil x+u\rfloor^{\alpha}<\left\lceil x-\frac{1}{2}x\right\rfloor^{\alpha}=2^{-\alpha}\lceil x\rfloor^{\alpha}.$
Multiplying by $x$ yields the result. Now, suppose $x>0$, which implies $-|u|>-\frac{1}{2}x$. Then we have
$\lceil x+u\rfloor^{\alpha}\geq\lceil x-|u|\rfloor^{\alpha}>\left\lceil x-\frac{1}{2}x\right\rfloor^{\alpha}=2^{-\alpha}\lceil x\rfloor^{\alpha}.$
We again multiply by $x$ to obtain the result.
With Lemma 3.4 at hand, we can differentiate $W$ along (19) to obtain the following for $|y|>2|u|$:
$\dot{W}$ $=-2y\lceil y-u_{1}\rfloor^{q_{1}}-2y\lceil y-u_{2}\rfloor^{q_{2}}+2yu_{1}u_{2}$
$\leq-2^{1-q_{1}}|y|^{q_{1}+1}-2^{1-q_{2}}|y|^{q_{2}+1}+2yu_{1}u_{2}$
$\leq-(2^{1-q_{1}}-2^{-q_{2}})|y|^{q_{1}+1}-2^{-q_{2}}|y|^{q_{2}+1}+2^{q_{2}}|u|^{4}$
$\leq-(2^{-q_{1}}-2^{-1-q_{2}})W^{\frac{1}{2}(q_{1}+1)}-2^{-1-q_{2}}W^{\frac{1}{2}(q_{2}+1)}$
$~{}~{}~{},\quad|y|>\max({\varrho}_{B}^{-1}(2^{q_{2}}|u|^{4}),2|u|),$
@@
$\varrho_{B}(r)=(2^{-q_{1}}-2^{-1-q_{2}})|r|^{q_{1}+1}+2^{-1-q_{2}}|r|^{q_{2}+1},$
which implies (19) is FxT ISS uniformly in $x$, and hence the assumptions are satisfied. To verify if (17) is FxT ISS, it remains to check if the interconnection terms of (17) satisfy the interconnection conditions (11). First, we define
$\tilde{V}$ $=V^{\frac{1}{4}(r_{1}+1)}+V^{\frac{1}{4}(r_{2}+1)}=|x|^{\frac{1}{2}(r_{1}+1)}+|x|^{\frac{1}{2}(r_{2}+1)}$
$\tilde{W}$ $=W^{\frac{1}{4}(q_{1}+1)}+W^{\frac{1}{4}(q_{2}+1)}=|y|^{\frac{1}{2}(q_{1}+1)}+|y|^{\frac{1}{2}(q_{2}+1)}.$
Then, we can compute $I_{1}$ and apply Lemmas 1-3 to obtain the estimate:
$\leq 2|x|\left(2^{1-r_{1}}|y|^{r_{1}}+K_{1-r_{2}}\left(|y|^{r_{2}}+|y||x|^{r_{2}-1}\right)\right)$
$\leq\dfrac{1}{\underline{\alpha}_{1}(c)}\left(|x|^{r_{1}+1}+|x|^{r_{2}+1}\right)+\overline{\alpha}_{1}(c)\left(|y|^{r_{1}+1}+|y|^{r_{2}+1}\right)$
$\leq\dfrac{1}{\underline{\alpha}_{1}(c)}\tilde{V}^{2}+2\overline{\alpha}_{1}(c)\tilde{W}^{2},$
for all $c>0$, where $\overline{\alpha}_{1},\underline{\alpha}_{1}\mathcal{\in}{K}_{\infty}$ are obtained by applying Lemma 1 twice and majorizing appropriately. We can pick $c>0$ sufficiently large such that (11a) and (11c) are satisfied, so it remains to check if $I_{2}$ satisfies (11b). Indeed, by computing $I_{2}$ we obtain:
$\leq 2|y|\left(|y|^{r_{1}}+|x|^{r_{1}}+2^{r_{2}-1}\left(|y|^{r_{2}}+|x|^{r_{2}}\right)+|u|\right)$
$\leq\dfrac{1}{\underline{\alpha}_{2}(c)}\left(|x|^{r_{1}+1}+|x|^{r_{2}+1}\right)+\overline{\alpha}_{2}(c)\left(|y|^{r_{1}+1}+|y|^{r_{2}+1}\right)$
$~{}~{}~{}+2|y|^{r_{1}+1}+2^{r_{2}}|y|^{r_{2}+1}+2|y||u|$
$\leq\dfrac{1}{\underline{\alpha}_{2}(c)}\tilde{V}^{2}+(2\overline{\alpha}_{2}(c)+3+2^{r_{2}})\tilde{W}^{2}+|u|^{2},$
for all $c>0$, where $\overline{\alpha}_{2},\underline{\alpha}_{2}\in\mathcal{K}_{\infty}$ are again obtained from Lemma 1. We conclude that the interconnection conditions are satisfied, and hence there exists $\varepsilon^{*}$ such that (17) is FxT ISS for $\varepsilon\in(0,\varepsilon^{*})$. We simulate system (17) with $r_{1}=\frac{1}{3},r_{2}=-\frac{1}{5},q_{1}=\frac{2}{5},q_{2}=-\frac{2}{7}$ and the following disturbances:
Figure 1: Trajectories of system (17) with (22), $[x_{0},z_{0}]=[3,2.37]$ and $\varepsilon=0.001$.
$u_{1}(t)$ $=\varepsilon_{0}e^{\sin{t}}$ (22a)
$u_{2}(t)$ $=\varepsilon_{0}\sin(19\log(t+1))-0.21.$ (22b)
We clearly have $u(t)=[u_{1}(t),u_{2}(t)]^{\top}\in\mathcal{L}_{\infty}^{2}$ and $u(t)<M\varepsilon_{0}$ for some $M>0$, so our computations indicate that the resulting system is FxT ISS with $\varepsilon_{0}$ treated as the “input”. The trajectories of this system are shown in Figure 1 for different values of $\varepsilon_{0}$.
4 Fixed-Time Feedback Optimization with Time-varying Costs
In this section, we leverage the results of Theorems 1-2 to study feedback optimization problems under slowly-varying cost functions. In particular, we consider the following class of optimization problems
$\min_{z\in\mathbb{R}^{n_{2}}}\phi_{\theta}(z),$ (23)
where $\theta$ is a time-varying parameter that evolves according to the following dynamics:
$\dot{\theta}=\varepsilon_{0}\Pi(\theta),\quad\theta\in\Theta.$ (24)
We let $z^{*}=\varphi(\theta)$ denote the solution to (23), where $\varphi(\cdot)$ is assumed to be $\mathcal{C}^{1}$ and unique.
We consider plants that satisfy Assumption 3:
$\dot{z}=g(x,z),$ (25)
In fact, for our purposes we do not need to consider (25) as having an input; we only require the boundary layer dynamics to be FxTS in $y$, uniformly in $x$. More formally, we impose the following assumption on (25):
@@ 4 Fixed-Time Feedback Optimization with Time-varying Costs
For (25), $g(x,z)=0$ if and only if $z=x$. Furthermore, there exists a $\mathcal{C}^{1}$ function $W:\mathbb{R}^{n_{2}}\to\mathbb{R}_{\geq 0}$ and $\beta_{1}>0,\beta_{2}\in\mathcal{K}_{\infty}$ such that
$\beta_{1}|y|^{2}\leq W(y)\leq\beta_{2}(|y|),$
and
$\dfrac{\partial W}{\partial y}g(x,y+x)\leq-\kappa_{1}W^{b_{1}}-\kappa_{2}W^{b_{2}},$
where $\kappa_{1},\kappa_{2}>0$, $b_{1}\in(0,1)$ and $b_{2}>1$. Moreover, there exists $\eta>0$ such that the function $W$ satisfies
$\left\lvert\dfrac{\partial W}{\partial y}\right\rvert$ $\leq\eta|y|,$ (26)
for all $y\in\mathbb{R}^{n_{2}}$. $\square$
For example, Assumption 4 is satisfied by the system
$\dot{z}=-\kappa_{1}\dfrac{z-x}{|z-x|^{2-2b_{1}}}-\kappa_{2}\dfrac{z-x}{|z-x|^{2-2b_{2}}},$
for some $\kappa_{1},\kappa_{2}>0$, $b_{1}\in(\frac{1}{2},1)$, $b_{2}>1$. In general, any system where $z$ converges to $x$ in fixed time for each $x$ will satisfy Assumption 4 if $W=|y|^{2}=|z-x|^{2}$ is a valid FxT Lyapunov function. Our goal is to design a control law on $x$ that can stabilize (25) in fixed-time while ensuring $z$ converges to the time-varying solution of (23). If measurements of $\nabla\phi_{\theta}(z)$ are available, we can consider the following update law for $x$:
$\dot{x}=-\varepsilon\left(\dfrac{\nabla\phi_{\theta}(z)}{|\nabla\phi_{\theta}(z)|^{\xi_{1}}}+\dfrac{\nabla\phi_{\theta}(z)}{|\nabla\phi_{\theta}(z)|^{\xi_{2}}}\right),$
where $\xi_{1}\in(0,1)$ and $\xi_{2}<0$. With the time-scale transformation $\tau=\varepsilon t$, we arrive at the following interconnection
$x^{\prime}$ $=-\dfrac{\nabla\phi_{\theta}(z)}{|\nabla\phi_{\theta}(z)|^{\xi_{1}}}-\dfrac{\nabla\phi_{\theta}(z)}{|\nabla\phi_{\theta}(z)|^{\xi_{2}}}$ (27a)
$\varepsilon z^{\prime}$ $=g(x,z)$ (27b)
where $x^{\prime}=\frac{dx}{d\tau}$ and $z^{\prime}=\frac{dz}{d\tau}$. Note that the stability properties of (27) can be recovered from the corresponding system with $\dot{x}$ and $\dot{z}$ instead of $x^{\prime}$ and $z^{\prime}$. Taking this into account, and applying the transformations $\tilde{x}=x-\varphi(\theta)$ and $\tilde{z}=z-\varphi(\theta)$, we arrive at the following system:
$\dot{\tilde{x}}$ $=-\dfrac{\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))}{|\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))|^{\xi_{1}}}-\dfrac{\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))}{|\nabla\phi_{\theta}(\tilde{z}+\varphi(\theta))|^{\xi_{2}}}-\varepsilon_{0}\dfrac{d\varphi(\theta)}{d\theta}\Pi(\theta)$ (28a)
$\varepsilon\dot{\tilde{z}}$ $=g(\tilde{x}+\varphi(\theta),\tilde{z}+\varphi(\theta))-\varepsilon\varepsilon_{0}\dfrac{d\varphi(\theta)}{d\theta}\Pi(\theta).$ (28b)
For this work, we will make the following assumptions on the structure of $\phi_{\theta}(z)$, which are fairly standard in the time-varying optimization literature:
Assumption 5
The functions $\phi_{\theta}$ satisfy the following:
  • There exists $L>0$ such that
    $|\nabla\phi_{\theta}(z_{1})-\nabla\phi_{\theta}(z_{2})|\leq L|z_{1}-z_{2}|,$ (29)
    for all $z_{1},z_{2}\in\mathbb{R}^{n_{2}}$ and $\theta\in\Theta$
  • There exists a $\mathcal{C}^{1}$ mapping $\varphi$ such that $z^{*}=\varphi(\theta)=\operatorname*{arg\,min}_{z}\phi_{\theta}(z)$, for each $\theta\in\Theta$.
  • There exists $\kappa>0$ such that $\phi_{\theta}(z_{2})\geq\phi_{\theta}(z_{1})+\nabla\phi_{\theta}(z_{1})^{\top}(z_{2}-z_{1})+\frac{\kappa}{2}|z_{2}-z_{1}|^{2}$, for all $z_{1},z_{2}\in\mathbb{R}^{n_{2}}$ and $\theta\in\Theta$.
$\square$
It is useful to note that Assumption 5 implies the following lemma:
Lemma 4.1.
If $\phi_{\theta}$ satisfies Assumption 5, the following hold:
$x^{\top}\nabla\phi_{\theta}(x+\varphi(\theta))$ $\geq\frac{1}{L}|\nabla\phi_{\theta}(x+\varphi(\theta))|^{2}$
$|\nabla\phi_{\theta}(x+\varphi(\theta))|$ $\geq\kappa|x|,$
for all $x\in\mathbb{R}^{n_{1}},\theta\in\Theta$. $\square$
We will show that (28) is FxT ISS with $\dot{\theta}$ being treated as the “input”, so slower variations of $\theta$ will result in a more accurate $\mathcal{GKL}$ bound on the flows of (28). Since we assume $\Theta$ is compact, meaning $\Pi(\theta)$ is bounded on $\Theta$, we can also view $\varepsilon_{0}$ as the “input”. For this interconnection we can state the following result:
Proposition 3.
Consider system (28) with $0<\xi_{1}<\min(2-2b_{1},1)$, $02-2b_{2}\leq\xi_{2}<$ 0 and $g$ satisfying Assumption 4. Furthermore, $\phi_{\theta}(z)$ satisfies Assumption 5 and $\theta$ evolves according to dynamics (24) where $\Theta$ is compact, $\Pi(\cdot)$ is Lipschitz and $\Theta$ is forward invariant under (24). Then, there exists $\varepsilon^{*}$ such that for all $\varepsilon\in(0,\varepsilon^{*})$, there exists $\beta_{\varepsilon}\in\mathcal{GKL}$ and $\vartheta_{\varepsilon}\in\mathcal{K}$ such that the following holds

### v2

@@ 3.3 A Stylized Example
$\dfrac{\partial V}{\partial x}f(x,h(x),u)$ $\leq-V^{{\tilde{r}_{1}}}{(x)}-V^{{\tilde{r}_{2}}}{(x)}+|u|^{2},$ (24)
where $\tilde{r}_{i}:=\frac{1}{2}(r_{i}+1)$. Thus, Assumption 2 is satisfied.
Similarly, we can differentiate $W$ along the trajectories of (23) to obtain the following:
$\dot{W}$ $\leq-(2^{-q_{1}}-2^{-1-q_{2}})W^{{\tilde{q}_{1}}}{(y)}-2^{-1-q_{2}}W^{{\tilde{q}_{2}}}{(y)},$
$~~~\quad\forall~|y|>\max\{{\varrho}_{B}^{-1}(2^{q_{2}}|u|^{4}),2|u|\},$
where $\tilde{q}_{i}=\frac{1}{2}(q_{i}+1)$ and
$\varrho_{B}(s):=(2^{-q_{1}}-2^{-1-q_{2}})|s|^{q_{1}+1}+2^{-1-q_{2}}|s|^{q_{2}+1}.$
This implies that system (23) is FxT ISS uniformly in $x$, and hence the assumptions are satisfied. To verify if system (21) is FxT ISS, it remains to check if the interconnection terms associated with system (21) satisfy the interconnection conditions (15). Let $\tilde{V}{(x)}=|x|^{{\tilde{r}_{1}}}+|x|^{{\tilde{r}_{2}}}$ and $\tilde{W}{(y)}=|y|^{{\tilde{q}_{1}}}+|y|^{{\tilde{q}_{2}}}$. Then, we can compute $I_{1}$ and apply Lemmas 1-2 to obtain the following estimate:
$\leq c\tilde{V}^{2}{(x)}+2c^{-\sigma_{1}}\tilde{W}^{2}{(y)},$
for all $c\in(0,1)$, where $\sigma_{1}>0$ is obtained by applying Lemma 1 twice and upper bounding appropriately. By taking $c>0$ sufficiently small we can satisfy conditions (15a) and (15c). Thus, it remains to check if $I_{2}$ also satisfies (15b). Indeed, by computing $I_{2}$ we obtain:
$\leq c\tilde{V}^{2}{(x)}+(2c^{-\sigma_{2}}+3+2^{r_{2}})\tilde{W}^{2}{(y)}+|u|^{2},$
for all $c\in(0,1)$, where $\sigma_{2}>0$ is again obtained from Lemma 1. We conclude that the conditions of Theorem 1 and 3 are satisfied, and hence there exists $\varepsilon^{*}$ such that the singularly perturbed system (21) is FxT ISS for $\varepsilon\in(0,\varepsilon^{*})$. We simulate the system using $r_{1}=\frac{2}{5},r_{2}=\frac{6}{5},q_{1}=\frac{1}{3},q_{2}=\frac{9}{7}$ and the disturbances $u_{1}(t)=e^{\sin{t}},u_{2}(t)=\sin(19\log(t+1))-0.21$, where we clearly have $u(t)=[u_{1}(t),u_{2}(t)]^{\top}\in\mathcal{L}_{\infty}^{2}$. The trajectories of this system with and without the input are shown in Figure 1, illustrating the FxT ISS property
Figure 1: Trajectories of system (21) with and without the disturbance $u(t)$, where we use $\varepsilon=0.01$. The theoretically computed settling time bound of $18.15$ obtained using (7) is conservative compared to the observed trajectories.
4 Fixed-Time Feedback Optimization with Time-varying Costs
In this section, we leverage the results of Theorems 1-3 to study a practical problem of interest in the context of singular perturbations: feedback optimization under slowly-varying cost functions [26]. In contrast to the existing asymptotic results [27, 28, 33, 34], we introduce an optimization-based controller able to achieve FxT stability via non-Lipschitz feedback. In particular, we consider plants of the form
$\dot{z}=g(\hat{x},z),$ (25)
where $z\in\mathbb{R}^{m}$ is the state, $\hat{x}\in\mathcal{L}_{\infty}^{n}$ is a measurable and bounded control input, and $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{m}$ is a continuous function satisfying the following condition:
Assumption 4
There exists a continuously differentiable, globally Lipschitz mapping $h:\mathbb{R}^{n}\to\mathbb{R}^{m}$ such that $g(\hat{x},h(\hat{x}))=0$ for all $\hat{x}\in\mathbb{R}^{n}$. Moreover, there exists a $\mathcal{C}^{1}$ function $W:\mathbb{R}^{m}\to\mathbb{R}_{\geq 0}$ and $c_{1},c_{2}>0$ such that
$c_{1}|y|^{2}$ $\leq W(y)\leq c_{2}|y|^{2},$ (26a)
$)\displaystyle\nabla W(y)^{\top}g(\hat{x},y+h(\hat{x}$)) $\leq-\kappa_{1}W^{b_{1}}{(y)}-\kappa_{2}W^{b_{2}}{(y)},$ (26b)
for all $y\in\mathbb{R}^{m}$, where $\kappa_{1},\kappa_{2}>0$, $b_{1}\in(0,1)$ and $b_{2}>1$. There also exists $\eta>0$ such that the function $W$ satisfies
$\left\lvert\nabla W(y)\right\rvert$ $\leq\eta|y|,$ (27)
for all $y\in\mathbb{R}^{m}$. $\square$
Assumption 4 is the “fixed-time” version of standard open-loop stability assumptions considered in the setting of feedback optimization [7, 27]. In particular, by taking y=zh(x^)y=z-h($\hat${x}) to quantify the deviation of $z$ from its steady-state approximation, the conditions of Assumption 4 simply ask that such deviation converges to zero by a fixed-time, for each fixed $\hat{x}\in\mathbb{R}^{n}$. For a general class of linear and nonlinear plants, this property can be achieved using different types of non-smooth controllers that combine super-linear and sub-linear feedback [35], homogeneity tools [36], or the implicit Lyapunov technique [37], to name just a few examples.
Our primary goal is to design a control law on the input $\hat{x}$ that stabilizes (25) in a fixed time the solution of the following time-varying optimization problem
$\min_{\hat{x},z}\ \phi_{\theta}(\hat{x},z)$ (28a)
[…]
where the time variation on the cost functions $\phi_{\theta}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is induced by a dynamic parameter $\theta\in\mathbb{R}^{q}$ that evolves according to the dynamics
$\dot{\theta}=\varepsilon\varepsilon_{0}\Pi(\theta),\quad\theta\in\Theta,$ (29)
where $\varepsilon_{0}\geq 0$, and $\varepsilon>0$ is a small parameter that captures the rate of change of $\theta$, and $\Pi:\mathbb{R}^{q}\to\mathbb{R}^{q}$ and $\Theta\subset\mathbb{R}^{q}$ satisfy the following mild conditions, which can be used to cover a broad class of time-varying signals $t\mapsto\theta(t):$
Assumption 5
The function $\Pi(\cdot)$ is Lipschitz continuous, and the set $\Theta$ is compact and forward invariant under the dynamics (29). $\square$
We can observe that, by substituting (28b) into (28a), we arrive at the unconstrained parameterized optimization problem:
[…]
where $)\Phi_{\theta}(\hat{x}):=\phi_{\theta}(\hat{x},h(\hat{x}$)). To guarantee that (30) is a well-defined optimization problem with a unique solution for each $\theta\in\mathbb{R}^{q}$, we consider cost functions that satisfy the following assumptions, which are fairly standard in the time-varying feedback optimization literature [33, 34]:
Assumption 6
[…]
Note that when $\varepsilon_{0}=0$ in (5), the parameter $\theta$ remains constant for all time, yielding a constant solution $\varphi(\theta)$ to (30). In contrast, when $\varepsilon_{0}\gg 0$, the function $t\mapsto\varphi(\theta(t))$ may exhibit fast time variations that are difficult to track without additional information about the functions $\Pi$, $h$, and $\phi_{\theta}$. Therefore, to address the optimization problem (28a) using real-time gradient feedback, we can regard $\varepsilon_{0}\Pi(\theta)$ as the “input” to the system and study FxT ISS with respect to the tracking error of $\varphi(\theta(\cdot))$.
4.1 Fixed-Time Gradient-Based Feedback
Given $\xi_{1}\in(0,1)$ and $\xi_{2}<0$, we define the following function $\mathcal{F}_{\xi_{1},\xi_{2}}:\mathbb{R}^{p}\to\mathbb{R}^{p}$:
$\mathcal{F}_{\xi_{1},\xi_{2}}(x)=\frac{x}{|x|^{\xi_{1}}}+\frac{x}{|x|^{\xi_{2}}},$ (31)
[…]
@@ 4.1 Fixed-Time Gradient-Based Feedback
$\hat{P}_{\theta}(\hat{x},z):=H(\hat{x})^{\top}\nabla\phi_{\theta}(\hat{x},z),\quad H(x)^{\top}=[\mathbb{I}_{n}\quad\mathbf{J}_{h}(x)^{\top}].$ (33)
It can be verified, using the chain rule, that
[…]
and hence, for each $\theta$, when the plant dynamics (25) are negligible, the dynamics (32) converge to the solution of (30) in fixed time [15, 16]. Since in our setting the dynamics (25) cannot be neglected, we can obtain a real-time feedback controller by replacing h(x^)h($\hat${x}) in (32) with the measured value of $z$ to obtain the following closed-loop system:
$\dot{z}$ $=g(\hat{x},z)$ (34a)
$\dot{\hat{x}}$ $=-\varepsilon\mathcal{F}_{\xi_{1},\xi_{2}}(\hat{P}_{\theta}(\hat{x},z)).$ (34b)
Figure 2: Trajectories of the FxT time-varying feedback optimization example, with $\varepsilon_{0}=5,0.2,0$ and varying initial conditions.
We study system (34) under the following mild Lipschitz assumption, which is standard in the literature [27]:
@@
For the function $\hat{P}_{\theta}(x,z)$ defined in (33), there exists $\ell>0$ such that
$|\hat{P}_{\theta}(\hat{x},z^{\prime})-\hat{P}_{\theta}(\hat{x},z)|\leq\ell|z^{\prime}-z|,$ (35)
for all $\hat{x}\in\mathbb{R}^{n}$, $z^{\prime},z\in\mathbb{R}^{m}$ and $\theta\in\mathbb{R}^{q}$. $\square$
To put system (34) into the form (1), let $x=\hat{x}-\varphi(\theta)$ and $\tau=\varepsilon t$, which leads to the following dynamics in the $\tau$-time scale:
$\varepsilon\frac{dz}{d\tau}$ $=g(x+\varphi(\theta),z)$ (36a)
$\frac{dx}{d\tau}$ $=-\mathcal{F}_{\xi_{1},\xi_{2}}(\hat{P}_{\theta}(x+\varphi(\theta),z))-\mathbf{J}_{\varphi}(\theta)u(\tau)$ (36b)
$\frac{d\theta}{d\tau}$ $=\varepsilon_{0}\Pi(\theta),\quad\theta\in\Theta,$ (36c)
where $u(\tau)=\varepsilon_{0}\Pi(\theta(\tau))$ can be thought of as the “input” in (36b). Note that, since $\Theta$ is compact and forward invariant, and the trajectories of (36c) are restricted to evolve in $\Theta$, we only need to consider the stability properties of system (36a)-(36b). The following result establishes FxT ISS for this system:
Theorem 4.
Suppose that Assumptions 4-6 hold, and consider the singularly perturbed system (36). Then, for all $\xi_{1}\in\left(0,\min(2-2b_{1},1)\right)$ and $\xi_{2}\in\left(2-2b_{2},0\right)$, there exists $\beta\in\mathcal{GKL}$, $\varrho\in\mathcal{K}$ and $\varepsilon^{*}>0$ such that